1. Explain the weighted average inventory cost method and give an example?
Answer #1

Weighted average inventory cost method is method of making weighted average of price for changing it to issue.

An example: ABC company has beginning balance of inventory of 500 units at $6 rate on 1st March; issued 445 units on March 14; purchase 250 units at $12 on 17th March; again 250 units of purchase at $14 on 25th March; and issued 260 units on 31st March.

Weighted average method

Date

Receive

Sales (Issue)

Balance

Qty.

Rate

Amount

Qty.

Rate

Amount

1

500

6

3,000

14

445

6

2,670

55

6

330

17

250 @12 = 3,000

55 + 250 = 305

3,330/305 = 10.92

330 + 3,000 = 3,330

25

250 @14 = 3,500

305 + 250 = 555

6830/555 = 12.31

3330+3500 = 6,830

31

260

12.31

3,200

555 – 260 = 295

3,630/295 = 12.31

6,830 – 3,200 = 3630

Ending inventory

3,630

Note: in the receive side all materials are purchased; materials are issued based on the most recent prices; the ending inventory at the month-end is $3,630.